OpenStudy (anonymous):

The fifth term of (a - 3b)5 is 405ab4 is this true or false ?

OpenStudy (anonymous):

true..

OpenStudy (anonymous):

The k+1 term of (a+b)^n is \[\left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k}b^k\]

OpenStudy (anonymous):

So the fifth term of (a-3b)^5 is\[\left(\begin{matrix}5 \\4\end{matrix}\right)a^{5-4}(-3b)^4=405ab^4\]

OpenStudy (anonymous):

Notice that the fifth term gets k value of four....

OpenStudy (anonymous):

thanckyou guys :)

OpenStudy (anonymous):

welcome